According to the quantity equation, if P = 6 and Y = 800, which of the...

  1. According to the quantity equation, if P = 6 and Y = 800, which of the following pairs could M and V be? (1 mark)
Answer #1

Answer

Option d

=====

quantity equation is:

PY=MV

P=price level

Y=real GDP

M=money supply

V=velocity

so the pair should have the product equal to PY

PY=6*800=$4800

working by options

MV=200*3=600

MV=400*4=1600

MV=600*5=3000

MV=800*6=4800

so the option d is correct as the pair equate the equation MV=PY=4800